OpenStudy (anonymous):

What is the value of summation (2i^2+i-8)? I=1

OpenStudy (anonymous):

\[{\Huge\sum_{i=1}^?}(2i^2+i-8)\] What is the upper limit?

OpenStudy (anonymous):

At any rate, these formulas should help: \[\sum_{i=1}^n1=n\\ \sum_{i=1}^ni=\frac{n(n+1)}{2}\\ \sum_{i=1}^ni^2=\frac{n(2n+1)(n+1)}{6}\\ \sum_{i=1}^n(ai+b)=a\sum_{i=1}^ni+b\sum_{i=1}^n1\]